# Hardware-inspired Continuous Variables Quantum Optical Neural Networks

### v1

@@ Appendix C Exact expression of a single-layer QONN with two-mode subtraction
with $\Re$ representing the real part. Now, let $\mathcal{D}_{\rho}:=\text{Tr}\big[\hat{a}_{q}^{\dagger}\hat{a}_{p}^{\dagger}\hat{a}_{p}\hat{a}_{q}\rho_{G}\big]$ denote the expression of the expectation value—trace—in the denominator, which represents the norm of $\rho$, and let $\mathcal{M}_{j|\rho}:=\text{Tr}\big[\hat{a}_{q}^{\dagger}\hat{a}_{p}^{\dagger}\hat{a}_{j}\hat{a}_{p}\hat{a}_{q}\rho_{G}\big]$ be the trace expression of the numerator, so
@@
The heralding denominator $D_{\rho}$ is the fourth-order Gaussian moment which, expanding about the means and applying Wick–Isserlis with the identities of Eq. (C1), gives the closed form
$\boxed{\;\begin{aligned} \mathcal{D_{\rho}}&=\big(|\bm{\bar{b}}_{p}|^{2}+\Sigma_{N+p,N+p}\big)\big(|\bm{\bar{b}}_{q}|^{2}+\Sigma_{N+q,N+q}\big)+\big|\Sigma_{N+p,N+q}\big|^{2}+\big|\Sigma_{p,N+q}\big|^{2}\\[-2.0pt] &\quad+\bm{\bar{b}}_{p}\,\bm{\bar{b}}_{q}^{\ast}\,\Sigma_{N+q,N+p}+\bm{\bar{b}}_{p}^{\ast}\,\bm{\bar{b}}_{q}\,\Sigma_{N+p,N+q}+\bm{\bar{b}}_{p}\,\bm{\bar{b}}_{q}\,\Sigma_{p,N+q}^{\ast}+\bm{\bar{b}}_{p}^{\ast}\,\bm{\bar{b}}_{q}^{\ast}\,\Sigma_{p,N+q}\,.\end{aligned}\;}$ (C5)
If the clicked subspace of $|G\rangle$ is diagonal ($\Sigma_{N+p,N+q}=\Sigma_{p,N+q}=0$), then $D_{\{p,q\}}=(|\bm{\bar{b}}_{p}|^{2}+\Sigma_{N+p,N+p})(|\bm{\bar{b}}_{q}|^{2}+\Sigma_{N+q,N+q})$.
@@
$\boxed{\ \begin{aligned} \mathcal{S}^{(2)}_{j|\rho}&=\Sigma_{j,p}\,\bm{\bar{b}}_{q}^{\ast}\,\bm{\bar{b}}_{p}\,\bm{\bar{b}}_{q}+\Sigma_{j,q}\,\bm{\bar{b}}_{p}^{\ast}\,\bm{\bar{b}}_{p}\,\bm{\bar{b}}_{q}\\ &\quad+\Sigma_{j,N+p}\,\bm{\bar{b}}_{q}^{\ast}\,\bm{\bar{b}}_{p}^{\ast}\,\bm{\bar{b}}_{q}+\Sigma_{j,N+q}\,\bm{\bar{b}}_{q}^{\ast}\,\bm{\bar{b}}_{p}^{\ast}\,\bm{\bar{b}}_{p}.\end{aligned}}$ (C7)
@@
Pick one mean among $\{q^{\dagger},p^{\dagger},p,q\}$ and pair the remaining four fluctuations by Wick (always including a pairing with $j$). The result is
$\boxed{\ \begin{aligned} \mathcal{S}^{(4)}_{j|\rho}&=\;\bm{\bar{b}}_{q}^{\ast}\Big[\Sigma_{j,p}\,\Sigma_{p,N+q}+\Sigma_{j,N+p}\,\Sigma_{N+p,N+q}+\Sigma_{j,N+q}\,\Sigma_{N+p,N+p}\Big]\\[2.0pt] &\quad+\bm{\bar{b}}_{p}^{\ast}\Big[\Sigma_{j,q}\,\Sigma_{p,N+q}+\Sigma_{j,N+p}\,\Sigma_{N+q,N+q}+\Sigma_{j,N+q}\,\Sigma_{N+q,N+p}\Big]\\[2.0pt] &\quad+\bm{\bar{b}}_{p}\Big[\Sigma_{j,q}\,\Sigma_{N+p,N+q}+\Sigma_{j,p}\,\Sigma_{N+q,N+q}+\Sigma_{j,N+q}\,\Sigma_{N+q,N+p}\Big]\\[2.0pt] &\quad+\bm{\bar{b}}_{q}\Big[\Sigma_{j,q}\,\Sigma_{N+p,N+p}+\Sigma_{j,p}\,\Sigma_{N+q,N+p}+\Sigma_{j,N+p}\,\Sigma_{N+q,N+p}\Big].\end{aligned}}$ (C8)
Thus, plugging Eqs. (C5), (C7) and (C8) into Eq. (C6), and integrating the result along with Eq. (C5) into Eq. (C4), one obtains the exact normalized output $\langle\hat{x}_{j}\rangle_{\rho}$ in closed form, expressed only in terms of $(\bm{\bm{\bar{b}}},\Sigma)$.
If the subtractions $\hat{A}_{K}$ are performed on modes that are not correlated after the squeezing operation (i.e. $\Sigma_{N+p,N+q}=\Sigma_{p,N+q}=0$), then $D_{\rho}$ factorizes and $\mathcal{S}^{(4)}_{j|\rho}$ loses all coupling terms, reducing the expression to a sum of single-mode activations. In the displacement-free limit where $\bm{\bar{b}}_{p}=\bm{\bar{b}}_{q}=0$ only those terms with entries of $\Sigma$ would survive. If $\Sigma_{r,N+s}=0$ for $r,s\in K$, then $\mathcal{S}^{(4)}_{j|pq}$ reduces to terms with $\Sigma_{N+r,N+s}$ only, which account for number covariances. Additionally, if $j\in K$, index coincidences cause some addends to combine, as well as in the case of $p=q$ where the terms would combine into a single ridge activation. However, the formulas above remain valid without modification for any of these cases.

### v2

@@ Appendix C Exact expression of a single-layer QONN with two-mode subtraction
with $\Re$ representing the real part. Now, let $\mathcal{D}_{\rho}:=\text{Tr}\big[\hat{a}_{q}^{\dagger}\hat{a}_{p}^{\dagger}\hat{a}_{p}\hat{a}_{q}\rho_{G}\big]$ denote the expression of the expectation value—trace—in the denominator, which represents the normalization factor of $\rho$ (and the subtraction success probability), and let $\mathcal{M}_{j|\rho}:=\text{Tr}\big[\hat{a}_{q}^{\dagger}\hat{a}_{p}^{\dagger}\hat{a}_{j}\hat{a}_{p}\hat{a}_{q}\rho_{G}\big]$ be the trace expression of the numerator, so
@@
The heralding denominator $\mathcal{D}_{\rho}$ is the fourth-order Gaussian moment which, expanding about the means and applying Wick–Isserlis with the identities of Eq. (C1), gives the closed form
$\boxed{\;\begin{aligned} \mathcal{D}_{\rho}&=\big(|\bm{\bar{b}}_{p}|^{2}+\Sigma_{N+p,N+p}\big)\big(|\bm{\bar{b}}_{q}|^{2}+\Sigma_{N+q,N+q}\big)+\big|\Sigma_{N+p,N+q}\big|^{2}+\big|\Sigma_{p,N+q}\big|^{2}\\[-2.0pt] &\quad+\bm{\bar{b}}_{p}\,\bm{\bar{b}}_{q}^{\ast}\,\Sigma_{N+q,N+p}+\bm{\bar{b}}_{p}^{\ast}\,\bm{\bar{b}}_{q}\,\Sigma_{N+p,N+q}+\bm{\bar{b}}_{p}\,\bm{\bar{b}}_{q}\,\Sigma_{N+q,p}+\bm{\bar{b}}_{p}^{\ast}\,\bm{\bar{b}}_{q}^{\ast}\,\Sigma_{p,N+q}\,.\end{aligned}\;}$ (C5)
@@
$\boxed{\ \begin{aligned} \mathcal{S}^{(2)}_{j|\rho}&=\bm{\bar{b}}_{q}^{\ast}\,\bm{\bar{b}}_{p}\,\bm{\bar{b}}_{q}\Sigma_{j,p}\,+\bm{\bar{b}}_{p}^{\ast}\,\bm{\bar{b}}_{p}\,\bm{\bar{b}}_{q}\Sigma_{j,q}\,+\bm{\bar{b}}_{q}^{\ast}\,\bm{\bar{b}}_{p}^{\ast}\,\bm{\bar{b}}_{q}\Sigma_{j,N+p}\,+\bm{\bar{b}}_{q}^{\ast}\,\bm{\bar{b}}_{p}^{\ast}\,\bm{\bar{b}}_{p}\Sigma_{j,N+q}\,.\end{aligned}}$ (C7)
@@
Pick one operator among $\{\hat{a}_{q}^{\dagger},\hat{a}_{p}^{\dagger},\hat{a}_{p},\hat{a}_{q}\}$ to contribute its mean, and pair the remaining four fluctuations into two-point correlators by Wick–Isserlis. The result is
$\boxed{\ \begin{aligned} \mathcal{S}^{(4)}_{j|\rho}=\ &\bm{\bar{b}}_{q}^{\ast}\Big[\Sigma_{j,p}\,\Sigma_{p,N+q}+\Sigma_{j,N+p}\,\Sigma_{N+q,N+p}+\Sigma_{j,N+q}\,\Sigma_{N+p,N+p}\Big]\\[2.0pt] +\ &\bm{\bar{b}}_{p}^{\ast}\Big[\Sigma_{j,q}\,\Sigma_{p,N+q}+\Sigma_{j,N+p}\,\Sigma_{N+q,N+q}+\Sigma_{j,N+q}\,\Sigma_{N+p,N+q}\Big]\\[2.0pt] +\ &\bm{\bar{b}}_{p}\Big[\Sigma_{j,q}\,\Sigma_{N+q,N+p}+\Sigma_{j,p}\,\Sigma_{N+q,N+q}+\Sigma_{j,N+q}\,\Sigma_{N+q,p}\Big]\\[2.0pt] +\ &\bm{\bar{b}}_{q}\Big[\Sigma_{j,q}\,\Sigma_{N+p,N+p}+\Sigma_{j,p}\,\Sigma_{N+p,N+q}+\Sigma_{j,N+p}\,\Sigma_{N+q,p}\Big].\end{aligned}}$ (C8)
Thus, substituting Eqs. (C5), (C7) and (C8) into Eq. (C6), and inserting the resulting $\mathcal{M}_{j|\rho}$ together with Eq. (C5) into Eq. (C4), yields the exact normalized output $\langle\hat{x}_{j}\rangle_{\rho}$ in closed form, expressed solely in terms of $\rho_{G}$’s statistics $(\bm{\bar{b}},\Sigma)$.
If the photon subtractions implemented by $\hat{A}_{K}=\hat{a}_{p}\hat{a}_{q}$ act on two modes that are mutually uncorrelated in $\rho_{G}$ (i.e., the inter-click correlations vanish, $\Sigma_{N+p,N+q}=\Sigma_{p,N+q}=0$), then the heralding denominator factorizes as $\mathcal{D}_{\rho}=(|\bm{\bar{b}}_{p}|^{2}+\Sigma_{N+p,N+p})(|\bm{\bar{b}}_{q}|^{2}+\Sigma_{N+q,N+q})$, and all terms in $\mathcal{S}^{(4)}_{j|\rho}$ that are proportional to cross-correlations between the clicked modes disappear. In this regime, the normalized readout of mode $j$ contains no inter-click coupling and can be interpreted as the sum of two single-mode (ridge) activations, one associated with mode $p$ and one with mode $q$. In the displacement-free limit for the clicked modes, $\bm{\bar{b}}_{p}=\bm{\bar{b}}_{q}=0$, one has $\mathcal{S}^{(2)}_{j|\rho}=\mathcal{S}^{(4)}_{j|\rho}=0$ and therefore $\langle\hat{x}_{j}\rangle_{\rho}=\sqrt{2}\,\Re(\bm{\bar{b}}_{j})$. More generally, if only the anomalous inter-click correlator vanishes ($\Sigma_{p,N+q}=0$ and, thus, $\Sigma_{N+q,p}=0$), then all addends proportional to such terms drop out, and any remaining inter-click dependence is mediated only through the normal correlator $\Sigma_{N+p,N+q}$ (and its Hermitian counterpart $\Sigma_{N+q,N+p}$), together with local variances and correlations involving the readout mode $j$. A single ridge activation is obtained either when one clicked mode is effectively irrelevant for the readout (e.g., $\bm{\bar{b}}_{q}=0$ and $\Sigma_{j,q}=\Sigma_{j,N+q}=0$, so only the $p$-activation remains), or in the degenerate case in which both subtraction events act on the same mode ($p=q$, so $\hat{A}=\hat{a}_{p}^{2}$), which collapses the response to a single (second-order) ridge activation on mode $p$; in this latter case, identical indices introduce combinatorial multiplicities and the corresponding expressions follow by explicitly re-evaluating the Wick pairings for $\hat{a}_{p}^{2}$. Finally, when $j\in\{p,q\}$, additional index coincidences cause some addends to combine accordingly.